OpenStudy (anonymous):

HELP 2 Math questions timed quiz!!!! Distribute: 4(3x + 7) (Points : 1) 7x + 7 12x + 7 7x + 28 12x + 28 OTHER ONE I AM TYPING BELOW

OpenStudy (anonymous):

\[Distribute: \frac{ 1 }{ 2 } (8x-16 )\]

OpenStudy (anonymous):

16x – 32 4x – 8 16x – 8 4x – 16

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

4(3x + 7) =12x + 28

OpenStudy (anonymous):

1/2(8x−16)=4x-8